OpenStudy (anonymous):

7^(x - 5) = 7^(4x + 7) I don't understand how to solve questions with expressions as exponents?

OpenStudy (jiteshmeghwal9):

If the bases are same at both sides then the powers are also equal so\[(x-5)=(4x+7)\]\[x-4x=7+5\]\[-3x=12\]\[x=\frac{12}{-3} \implies -4\]

OpenStudy (anonymous):

If x^a = x^b, then a = b

OpenStudy (anonymous):

Oh, I see. What if the bases are not the same?

OpenStudy (jiteshmeghwal9):

then try to make the bases same:)

OpenStudy (jiteshmeghwal9):

like\[2^x=4\]\[2^x=2^2\]\[x=2\]

OpenStudy (anonymous):

Oh, alright I think I get it now. Thanks :)